OpenStudy (across):

I was looking through my old tests and found a short but interesting problem. You guys give it a quick try! --- A culture initially has \(N_0\) number of bacteria. At time \(t=1\) hour the number of bacteria is measured to be \((3/2)N_0\). If the rate of growth is proportional to the number of bacteria present, determine the time necessary for the number of bacteria to triple. Hint: the model to use is\[\frac{dP}{dt}=P.\]

OpenStudy (across):

My apologies, the model should be\[\frac{dP}{dt}=kP,\]where \(k\) is a constant. It has to be more general.

OpenStudy (dumbcow):

exponential growth problem: P = e^kt Let N0 = 1 3/2 = e^k --> k = ln(3/2) Plug in to find t when P = 3N0 e^kt = 3 (3/2)^t = 3 --> t = ln(3)/ln(3/2) = 2.71

OpenStudy (across):

@dumbcow: you got it! :P

OpenStudy (anonymous):

If the number of bacteria at time t is \(P(t)\), then the rate of change is given by: \[\frac{dP}{dt}=cP\] The solution of this Homogeneous DE is \(P(t)=ke^{ct}.\) Then at time t=0 \(P(0)=k=N_0\) and at t=1, \(\frac{3}{2}N_0=N_0e^c \implies c=\ln{\frac{3}{2}}.\)

OpenStudy (anonymous):

Now, we have the formula we can find thethe number of bacteria at any given time t. Or we can find when it's troubled too.

OpenStudy (pokemon23):

what kind of math is this?

OpenStudy (across):

Ordinary differential equations.

OpenStudy (pokemon23):

in English please

OpenStudy (pokemon23):

:0

OpenStudy (across):

It's a post-calculus subject. :)

OpenStudy (anonymous):

pokemon!! How's it going?

OpenStudy (anonymous):

@pokemon23: Something we take after a 1-year of single variable calculus.

OpenStudy (pokemon23):

hey AnwarA going fine :) I finally understand the concept of electrical circuits

OpenStudy (anonymous):

pokemon's head: "What's single-variable calculus?" :P

OpenStudy (across):

^ LOL

OpenStudy (pokemon23):

ok thanks ffm I'm in Algebra hopefully I will go into further into math because I want to be an engineer

OpenStudy (pokemon23):

I don't how to answer that question

OpenStudy (anonymous):

some people take this simultaneously with multi-variable and some even with Linear algebra.

OpenStudy (anonymous):

Good luck pokemon! :D

OpenStudy (katrinakaif):

Why are you twisting with electrical circuits when you are taking Algebra ?

OpenStudy (pokemon23):

I take electricity in hs (freshman)

OpenStudy (across):

@FoolForMath: That's right! I was then taking cal 3, and most of my friends were taking linear algebra.

OpenStudy (anonymous):

Linear algebra it's very helpful with vec calc

OpenStudy (pokemon23):

math these days

OpenStudy (anonymous):

You want to become an Electrical Engineer? @pokemon

OpenStudy (katrinakaif):

Math is more than helpful. Its even used to figure crimes

OpenStudy (anonymous):

Linear algebra is very useful to multi-variable too, lets say we can't understand multivariable calculus without moderate understanding of Linear algebra.

OpenStudy (anonymous):

@Katrina like the ones of Numb3rs?

OpenStudy (across):

@KatrinaKaif: Numb3rs was a cool show!

OpenStudy (pokemon23):

well yes anwar also become a pilot in aviation

OpenStudy (anonymous):

And a doctor too?

OpenStudy (anonymous):

Maybe a lawyer as well? ;P

OpenStudy (anonymous):

@across and @katrina: I love Numb3rs too, Markov chain and max-min network flows rules :D

OpenStudy (pokemon23):

nah I can't be a doctor

OpenStudy (pokemon23):

I hate looking at blood

OpenStudy (katrinakaif):

Most definitively! Its absurd that such mechanics in math comes useful. They do have mathematicians working on crime!

OpenStudy (pokemon23):

freaks me out

OpenStudy (anonymous):

I finally found Numb3rs fans haha

OpenStudy (anonymous):

Haha, just teasing you man! I studied Electrical Engineering for 3 semesters, I can probably help if you want to ask anything! :D @pokemon

OpenStudy (anonymous):

Mathematics is great tool, however I don't know how people study pure mathematics, it's the application that make it so AWESOME :D

OpenStudy (anonymous):

Hi got a question!

OpenStudy (anonymous):

I sorry

OpenStudy (anonymous):

across is the best one to answer you FFM!

OpenStudy (anonymous):

@AnwarA: Liar.. you told me that you are a pure math student :P

OpenStudy (katrinakaif):

I'd like a criminal investigating job =)

OpenStudy (across):

@FoolForMath: I like to think of it this way: Pure mathematicians develop the tools whilst applied mathematicians use them!

OpenStudy (anonymous):

I didn't!

OpenStudy (anonymous):

Yeah that too applied kat :D

OpenStudy (katrinakaif):

Sheikhs are liars. Processed and Confirmed.

OpenStudy (pokemon23):

thanks AnwarA I will gladly need your help when I get a midterm in electricity circuits

OpenStudy (anonymous):

lol @kat

OpenStudy (anonymous):

What did I do to you Katrina? :(

OpenStudy (anonymous):

I don't know where the Kirchoff laws come from

OpenStudy (anonymous):

I think across would kick us if she can! I'm off to chat! :)

OpenStudy (anonymous):

@across: Yes I agree, but what's the fun in developing when you are not applying it ? ;)

OpenStudy (anonymous):

well I know they come from Maxwell equations, but i cand not derived it by myself

OpenStudy (katrinakaif):

Lol Anwar =) Just my way of communicating to you

OpenStudy (anonymous):

@victorarana: Let's consider something more elementary, I know Fermat's last theorem but I can't prove it :P

OpenStudy (across):

@FFM: You get to brag about the usefulness of your tools! No, but I know what you mean. :P I simply love the elegance of it.

OpenStudy (anonymous):

I don't want a proof I just want to see the process of derivation.

OpenStudy (anonymous):

Well the theorem is elementary but not the proofs :P

OpenStudy (katrinakaif):

Derivatives are more than easy =P

OpenStudy (anonymous):

@KatrinaKaif: Well not in the time of Newton ;)

OpenStudy (across):

They're easy right now because there were guys who cracked their heads proving\[f(x)=x^n,\]\[f'(x)=nx^{n-1}\](though that one's not hard to derive).

OpenStudy (across):

;P

OpenStudy (anonymous):

LOL!

OpenStudy (anonymous):

@across: 100 years from now people will say Measure theory is easy :P

OpenStudy (anonymous):

Will they?

OpenStudy (across):

I completely agree with you. lol Also, calculus will be an elementary school course.